# Speculative Ensemble: Fast Large Language Model Ensemble via Speculation

### v1

@@ 3.3 Alternate Proposal Framework
In  Section 3.2, we explore the application of speculative decoding to LLM ensembles. However, we don’t consider the bonus token, that is, the additional token generated when all proposal tokens are accepted. This is because the bonus token follows the distribution of the verifier rather than the ensemble distribution and can not be directly appended to the output sequence. In this subsection, we introduce an ensemble framework, termed the alternate proposal framework, which effectively leverages the bonus token and demonstrates superior performance.
As shown in Figure 2, in the alternate proposal framework, the generation of a bonus token is treated as a proposal from the current verifier, which is subsequently verified by the current proposer. Specifically, let the proposer be denoted as $\mathcal{M}_{q}$ and the verifier as $\mathcal{M}_{p}$ with proposal lengths $\gamma_{q}$ and $\gamma_{p}$, respectively. If all tokens proposed by $\mathcal{M}_{q}$ are accepted, a total of $\gamma_{q}+1$ tokens will be generated. The first $\gamma_{q}$ tokens follows the distribution $r_{i+j}(x)=\mathcal{E}(q_{i+j}(x),p_{i+j}(x))$, for $j=1,\dots,\gamma_{q}$, while the $\gamma_{q}+1$-th token is drawn from $p_{i+\gamma_{q}+1}(x)$. At this stage, the $\gamma_{q}+1$-th token, referred to as bonus token, is treated as the initial token in $\mathcal{M}_{p}$’s proposal. Subsequently, $\mathcal{M}_{p}$ will generate an additional $\gamma_{p}-1$ tokens to complete its proposal.
@@
In Corollary 3.5, we proved that in the weighted ensemble scenario, the speculative-based ensemble is never slower than the vanilla ensemble and is typically faster. In this subsection, with the alternate proposal framework, we extend this conclusion to other types of two-model ensembles.
@@
For any two models, there exist values of $\gamma_{q}$ and $\gamma_{p}$ such that the speed of the alternate proposal framework is never slower than the vanilla ensemble and is almost always faster.
@@
Consider $\gamma_{q}=\gamma_{p}=1$, then $\frac{(1-\alpha^{\gamma_{q}})(1+c)}{(1-\alpha)(1+c\gamma_{q}-\alpha^{\gamma_{p}}c)}\geq 1$ holds universally. The equality holds only when $c=0$ or $\alpha=0$. However, $c>0$ because the execution time of the proposal model is non-negligible, and $\alpha>0$ holds unless the proposal distribution and the ensemble distribution do not overlap, which is almost impossible. ∎
An intuitive interpretation of Corollary 3.7 is that when $\gamma_{q}=\gamma_{p}=1$, even in the worst-case scenario—where all proposal tokens are rejected—each token generation still requires only one proposal and one verification. This results in the same number of model invocations as the standard ensemble. In practice, however, it is rare for all proposal tokens to be rejected. Once a token is accepted, the ensemble process becomes more efficient.
@@ 3.4 Generalize to More Models
Figure 3: The sketch of SE in three-model ensemble scenario. The colored boxes represent the stored probability distributions, while the grey boxes represent the discarded ones. Each invocation involves scoring the current proposal tokens and generating a bonus token. For clarity, we assume that the proposal length for each model is 1 and that all proposed tokens are accepted.
In this subsection, we extend SE to the $n$-model ensemble scenario. The core principles remain similar to the two-model case, with acceleration driven by two key factors. First, each model can score the proposals of other models in parallel, where scoring refers to computing the probability distribution of a proposal from other models. [^1] We use the term ”scoring” rather than ”verification” because, unlike in the two-model case, scoring does not immediately trigger verification; instead, verification occurs only after all models have scored a token. Second, during scoring, a model can naturally generate a bonus token, which further improves efficiency. We illustrate the speculative ensemble process in the $n$-model scenario with a simple example, while detailed pseudocode and a general visualization are provided in Section B.2.
As shown in Figure 3, the process begins in step 1 with the default proposal model, $\mathcal{M}_{1}$, generating a proposal token $x_{1}$. In step 2, $\mathcal{M}_{2}$ scores $x_{1}$ while simultaneously generating a bonus token $x_{2}$. Similarly, in step 3, $\mathcal{M}_{3}$ scores both $x_{1}$ and $x_{2}$ in parallel and produces another bonus token, $x_{3}$. At this point, $x_{1}$ has been scored by both $\mathcal{M}_{2}$ and $\mathcal{M}_{3}$, enabling the computation of its ensemble distribution $r_{1}(x)$ for verification. The associated distributions $p_{1}^{(1)}(x)$, $p_{2}^{(1)}(x)$, $p_{3}^{(1)}(x)$ are no longer needed and are discarded.
@@ 4.1 Experimental Setups
Datasets and evaluation. We test SE across multiple tasks including code generation, mathematical reasoning, multi-task understanding, and text summarization on HumanEval (Chen et al., 2021), GSM8K (Cobbe et al., 2021), MMLU (Hendrycks et al., 2021) and CNNDM (See et al., 2017), respectively. We measure each method’s speed by the average tokens generated per second and compute the speedup ratio relative to the standard ensemble.
Ensemble functions and methods. We experiment with two ensemble functions: weighted ensemble (WE) at the distribution level (Equation 4) and contrastive decoding (CD) at the logits level (Equation 5). For WE, in the two-model case, we set $\lambda=0.5$ and temperature $T=1$; in the three-model case, each model’s coefficient was set to $1/3$. For CD, we set $\mu=0.1$, which is the most common setting, and set $T$ to both 0 and 1. WE with $T=0$ is not tested due to its uncommon use, as it leads to a one-hot distribution, reducing information. Among two ensemble functions, three methods are compared: (1) the standard ensemble (WE, CD); (2) an accelerated version with speculative decoding (SD), using the smallest model as the proposal and the ensemble model as the target (WE-SD, CD-SD); and (3) Speculative Ensemble (WE-SE, CD-SE).
Model pair configuration. We experiment on different types of LLMs, including Llama-2 (Touvron et al., 2023; Miao et al., 2024), Vicuna (Zheng et al., 2023), Llama-3 (Dubey et al., 2024), Qwen-2.5 (Team, 2024), and OPT (Zhang et al., 2022). Model pair configurations for each ensemble are in Table 1. We also test a three-model ensemble using Qwen2.5-1.5B-Instruct and its code and math versions in the WE setting.
@@
WE Llama-Vicuna Llama-2-7B Vicuna-7B-V1.5
@@
CD Llama-3 Llama-3.2-1B Llama-3.1-8B-Instruct
@@
Configuration of $\gamma$. The proposal length $\gamma$ is the only hyperparameter in SD, affecting the algorithm’s acceleration. In the two-model SE setting, $\gamma$ corresponds to the proposal length of the smaller model, with the larger model fixed at 1. For simplicity, we refer to the smaller model with $\gamma>1$ as the proposal model of SE, since it typically serves this role. We tested $\gamma=5$ and $\gamma=1$ for SE and SD speeds, reporting the optimal results. $\gamma=5$ is the common setting, while $\gamma=1$ ensures acceleration (Corollary 3.7). In the three-model SE, all models have a proposal length of 1.
@@ 4.2 Main Results
Table 2 and Table 3 display the speedup ratios for each method relative to the standard ensemble in the WE and CD settings, respectively. [^2] The results of OPT model pair are shown in Section C.2. From these two tables, we have the following findings. First, SE not only consistently achieves the highest speedup in all settings, it also gets speedup across all settings, which supports the findings in Corollary 3.7. In contrast, SD may reduce the ensemble speed in some cases. For example, when using the Llama-2 model pair with $T=1$ on HumanEval in Table 3, applying SD reduces the speed to 0.94x of the standard ensemble. A similar speed reduction was also observed in the three-model scenario in Table 2. This is because vanilla SD does not inherently ensure acceleration. When the acceptance rate is low, SD may perform slower than standard decoding.
Second, compared to the CD scenario, the WE scenario ensures a higher minimum speedup for SE. In the two-model case, SE achieves a minimum speedup of 1.34x, while in the three-model case, it reaches at least 1.27x. In contrast, the CD scenario has a speedup as low as 1.11x. This difference arises because SE maintains a consistently high acceptance rate in the WE scenario, as outlined in Corollary 3.3.
Third, the speedup varies across tasks and is influenced by the determinism of task outputs. For example, in the WE scenario, SE achieved the highest speedup on HumanEval, averaging 1.65x, as code generation demands strictly formatted outputs. Conversely, SE has a lower speedup of 1.36x on a text summarization task, where output flexibility is higher. This difference stems from the alignment between the proposal and target models: in highly deterministic tasks, their outputs exhibit greater similarity, leading to a higher acceptance rate and, consequently, stronger acceleration.
@@ 4.3 Analysis
Speedup ratio for different weight $\lambda$ in WE. We examine the speedup effect of SE when $\lambda$ takes values other than just 0.5. Specifically, we conduct experiments with $\lambda$ values ranging from 0.1 to 0.9, using the Llama-Vicuna and Qwen-3b model pairs on the HumanEval and GSM8K datasets. The results, presented in Figure 4, show that SE consistently achieves a high speedup of at least 1.5x across all tested $\lambda$ values. This consistent speedup occurs because, when the two models are of similar sizes, for any $\lambda$, an appropriate proposal model can be selected to maintain a high acceptance rate during SE process (as explained in Corollary 3.3), ensuring the observed speedup.
Speedup ratio for different weight values of $\mu$ in CD. Similarly, we examine the speedup effect of CD when $\mu$ takes other values. Specifically, we conduct experiments with $\mu$ ranging from 0.1 to 0.5, using the Llama-3 model pair on the HumanEval and GSM8K datasets. The results, presented in Figure 5, show that SE consistently accelerates the CD process across all tested $\mu$.
Furthermore, we find that an increase in $\mu$ results in a reduced speedup. This occurs because CD computes the ensemble distribution by subtracting the proposal model’s information from the target model’s. As $\mu$ grows, the gap between these distributions widens, lowering the acceptance rate (Equation 7). Despite this, the speedup ratio remains above 1.00x, confirming that CD-SE always accelerates, consistent with Corollary 3.7.
@@ 5 Conclusion
This paper introduces Speculative Ensemble (SE), an extension of speculative decoding that accelerates ensemble inference while maintaining output quality. SE refines the verification mechanism for direct ensemble sampling and introduces an alternate proposal framework to further boost efficiency. We demonstrate the effectiveness of SE through both theoretical analysis and empirical validation.
@@ Appendix A Mathematical Proofs
A.1 Correctness of Speculative Ensemble
Assume that in speculative decoding, the proposal distribution is $q(x)$, the target distribution is $p(x)$, and the ensemble distribution is $r(x)=\mathcal{E}(q(x),p(x))$. Referring to the proof of speculative decoding correctness (Leviathan et al., 2023), we now prove that the sampling method described in  Section 3.2 ensures that the generated tokens align with the ensemble distribution $r(x)$.
@@ A.1 Correctness of Speculative Ensemble
Assume that the time required to invoke the target model once is $T$, and the time required to invoke a proposal model is $cT$. Therefore, one proposal and one verification together take $(\gamma c+1)T$ time. The time required to generate one token is $\frac{(1-\alpha)(\gamma c+1)}{1-\alpha^{\gamma}}T$. In the vanilla ensemble, the time required to generate one token is $(c+1)T$. Therefore, the improvement factor in total walltime is $\frac{(1-\alpha^{\gamma})(1+c)}{(1-\alpha)(\gamma c+1)}$.
@@ B.1 Alternate Proposal Framework
Algorithm 1 The alternate proposal framework. Propose takes the current token sequence $S$ and a constant $\gamma$ as inputs and generates $\gamma$ tokens $T$. Score feeds a sequence to the current verifier to obtain logits $L$ and a bonus token $t$. Verify examines $T$ to decide whether it should be accepted according to ensemble logits.
@@
     $L^{\prime}=\mathcal{E}^{\prime}(L)$
@@
B.2 Speculative Ensemble Framework
The speculative ensemble framework is a generalization of the alternate proposal framework to scenarios involving more than three models. Figure 7 illustrates our speculative ensemble in a three-model scenario, where the models $\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}$ have proposal lengths of $\gamma_{1}=3$, $\gamma_{2}=2$, and $\gamma_{3}=1$, respectively.

### v2

@@ 3.3 Alternate Proposal Framework
In  Section 3.2, we explore the application of speculative decoding to LLM collaboration. However, we don’t consider the bonus token, that is, the additional token generated when all proposal tokens are accepted. This is because the bonus token follows the distribution of the verifier rather than the combined distribution and can not be directly appended to the output sequence. In this subsection, we introduce a collaboration framework, termed the alternate proposal framework, which effectively leverages the bonus token and demonstrates superior performance.
As shown in Figure 2, in the alternate proposal framework, the generation of a bonus token is treated as a proposal from the current verifier, which is subsequently verified by the current proposer. Specifically, let the proposer be denoted as $\mathcal{M}_{q}$ and the verifier as $\mathcal{M}_{p}$ with proposal lengths $\gamma_{q}$ and $\gamma_{p}$, respectively. If all tokens proposed by $\mathcal{M}_{q}$ are accepted, a total of $\gamma_{q}+1$ tokens will be generated. The first $\gamma_{q}$ tokens follows the distribution $r_{i+j}(x)=\mathcal{C}(q_{i+j}(x),p_{i+j}(x))$, for $j=1,\dots,\gamma_{q}$, while the $\gamma_{q}+1$-th token is drawn from $p_{i+\gamma_{q}+1}(x)$. At this stage, the $\gamma_{q}+1$-th token, referred to as bonus token, is treated as the initial token in $\mathcal{M}_{p}$’s proposal. Subsequently, $\mathcal{M}_{p}$ will generate an additional $\gamma_{p}-1$ tokens to complete its proposal.
@@
In Corollary 3.5, we proved that in the weighted ensemble scenario, the Naive-CoS is never slower than the vanilla collaborative decoding and is typically faster. In this subsection, with the alternate proposal framework, we extend this conclusion to any form of two-model collaboration.
@@
For any two models, there exist values of $\gamma_{q}$ and $\gamma_{p}$ such that the speed of the alternate proposal framework is never slower than the vanilla collaboration and is almost always faster.
@@
Consider $\gamma_{q}=\gamma_{p}=1$, then $\frac{(1-\alpha^{\gamma_{q}})(1+c)}{(1-\alpha)(1+c\gamma_{q}-\alpha^{\gamma_{p}}c)}\geq 1$ holds universally. The equality holds only when $c=0$ or $\alpha=0$. However, $c>0$ because the execution time of the proposal model is non-negligible, and $\alpha>0$ holds unless the proposal distribution and the combined distribution do not overlap, which is almost impossible. ∎
An intuitive interpretation of Corollary 3.7 is that when $\gamma_{q}=\gamma_{p}=1$, even in the worst-case scenario—where all proposal tokens are rejected—each token generation still requires only one proposal and one verification. This results in the same number of model invocations as the standard collaborative decoding. In practice, however, it is rare for all proposal tokens to be rejected. Once a token is accepted, the collaboration process becomes more efficient.
@@ 3.4 Generalize to More Models
Figure 3: The sketch of CoS in three-model collaboration scenario. The colored boxes represent the stored probability distributions, while the gray boxes represent the discarded ones. Each invocation involves scoring the current proposal tokens and generating a bonus token. For clarity, we assume that the proposal length for each model is 1 and that all proposed tokens are accepted.
In this subsection, we extend CoS to the $n$-model collaboration scenario. The core principles remain similar to the two-model case, with acceleration driven by two key factors. First, each model can score the proposals of other models in parallel, where scoring refers to computing the probability distribution of a proposal from other models. [^1] We use the term “scoring” rather than “verification” because, unlike in the two-model case, scoring does not immediately trigger verification; instead, verification occurs only after all models have scored a token. Second, during scoring, a model can naturally generate a bonus token, which further improves efficiency. We illustrate the CoS process in the $n$-model scenario with a simple example, while detailed pseudocode and a general visualization are provided in Section B.2.
As shown in Figure 3, the process begins in step 1 with the default proposal model, $\mathcal{M}_{1}$, generating a proposal token $x_{1}$. In step 2, $\mathcal{M}_{2}$ scores $x_{1}$ while simultaneously generating a bonus token $x_{2}$. Similarly, in step 3, $\mathcal{M}_{3}$ scores both $x_{1}$ and $x_{2}$ in parallel and produces another bonus token, $x_{3}$. At this point, $x_{1}$ has been scored by both $\mathcal{M}_{2}$ and $\mathcal{M}_{3}$, enabling the computation of its combined distribution $r_{1}(x)$ for verification. The associated distributions $p_{1}^{(1)}(x)$, $p_{2}^{(1)}(x)$, $p_{3}^{(1)}(x)$ are no longer needed and are discarded.
@@ 4.1 Experimental Setups
Datasets and evaluation. We test CoS across multiple tasks including code generation, mathematical reasoning, multi-task understanding, and text summarization on HumanEval (Chen et al., 2021), GSM8K (Cobbe et al., 2021), MMLU (Hendrycks et al., 2021), and CNNDM (See et al., 2017), respectively. We measure each method’s speed by the average tokens generated per second and compute the speedup ratio relative to the standard collaborative decoding. All experiments are conducted on RTX 3090, except for evaluations involving the Llama-Vicuna model pair, which use the A6000 GPU. Additionally, we also test on the Ascend 910B3 NPU; the corresponding results are shown in Table 9 and Table 10.
Combination functions and methods. We experiment with two combination functions: weighted ensemble (WE) at the distribution level (Equation 4) and contrastive decoding (CD) at the logits level (Equation 5). For WE, in the two-model case, we set $\lambda=0.5$ and temperature $T=1$; in the three-model case, each model’s coefficient was set to $1/3$. For CD, we set $\mu=0.1$, which is the most common setting, and set $T$ to both 0 and 1. WE with $T=0$ is not tested due to its uncommon use, as it leads to a one-hot distribution, reducing information. Among two combination functions, four methods are compared: (1) the standard collaborative decoding (WE, CD); (1) parallel collaborative decoding (WE-P, CD-P); (2) an accelerated version with speculative decoding (SD), using the smallest model as the proposal and the combined distribution as the target (WE-SD, CD-SD); and (3) CoS (WE-CoS, CD-CoS). Since SCD is equivalent to Naive-CoS, its results are included in our ablation on alternative proposal frameworks (Section C.4).
Model pair configuration. We experiment on different types of LLMs, including Llama-2 (Touvron et al., 2023; Miao et al., 2024), Vicuna (Zheng et al., 2023), Llama-3 (Dubey et al., 2024), Qwen-2.5 (Team, 2024), and OPT (Zhang et al., 2022). Model pair configurations for each combination function are in Table 1. We also test a three-model collaboration using Qwen2.5-1.5B-Instruct and its code and math versions in the WE setting.
@@
Weight Ensemble (WE)
Llama-Vicuna Llama-2-7B Vicuna-7B-V1.5
@@
Contrastive Decoding (CD)
Llama-3 Llama-3.2-1B Llama-3.1-8B-Instruct
@@
Configuration of $\gamma$. The proposal length $\gamma$ is the only hyperparameter in SD, affecting the algorithm’s acceleration. In the two-model CoS setting, $\gamma$ corresponds to the proposal length of the smaller model, with the larger model fixed at 1. For simplicity, we refer to the smaller model with $\gamma>1$ as the proposal model of CoS, since it typically serves this role. We tested $\gamma=5$ and $\gamma=1$ for CoS and SD speeds, reporting the optimal results. $\gamma=5$ is the common setting, while $\gamma=1$ ensures acceleration (Corollary 3.7). In the three-model CoS, all models have a proposal length of 1.
@@ 4.2 Main Results
Table 2 and Table 3 display the speedup ratios for each method relative to the standard collaborative decoding in the WE and CD settings, respectively. [^2] The results of OPT model pair are shown in Section C.2. From these two tables, we have the following findings. First, CoS not only consistently achieves the highest speedup in all settings, it also gets speedup across all settings, which supports the findings in Corollary 3.7. In contrast, SD may reduce the collaboration speed in some cases. For example, when using the Llama-2 model pair with $T=1$ on HumanEval in Table 3, applying SD reduces the speed to 0.94x of the standard collaboration. A similar speed reduction was also observed in the three-model scenario in Table 2. This is because vanilla SD does not inherently ensure acceleration. When the acceptance rate is low, SD may perform slower than standard decoding.
Second, compared to the CD scenario, the WE scenario ensures a higher minimum speedup for CoS. In the two-model case, CoS achieves a minimum speedup of 1.34x, while in the three-model case, it reaches at least 1.27x. In contrast, the CD scenario has a speedup as low as 1.11x. This difference arises because CoS maintains a consistently high acceptance rate in the WE scenario, as outlined in Corollary 3.3.
Third, the speedup varies across tasks and is influenced by the determinism of task outputs. For example, in the WE scenario, CoS achieved the highest speedup on HumanEval, averaging 1.65x, as code generation demands strictly formatted outputs. Conversely, CoS has a lower speedup of 1.36x on a text summarization task, where output flexibility is higher. This difference stems from the alignment between the proposal and target models: in highly deterministic tasks, their outputs exhibit greater similarity, leading to a higher acceptance rate and, consequently, stronger acceleration.
@@ 4.3 Analysis
Speedup ratio for different weight $\lambda$ in WE. We examine the speedup effect of CoS when $\lambda$ takes values other than just 0.5. Specifically, we conduct experiments with $\lambda$ values ranging from 0.1 to 0.9, using the Llama-Vicuna and Qwen-3b model pairs on the HumanEval and GSM8K datasets. The results, presented in Figure 4, show that CoS consistently achieves a high speedup of at least 1.5x across all tested $\lambda$ values. This consistent speedup occurs because, when the two models are of similar sizes, for any $\lambda$, an appropriate proposal model can be selected to maintain a high acceptance rate during CoS process (as explained in Corollary 3.3), ensuring the observed speedup.
Speedup ratio for different weight values of $\mu$ in CD. Similarly, we examine the speedup effect of CD when $\mu$ takes other values. Specifically, we conduct experiments with $\mu$ ranging from 0.1 to 0.5, using the Llama-3 model pair on the HumanEval and GSM8K datasets. The results, presented in Figure 5, show that CoS consistently accelerates the CD process across all tested $\mu$.
Furthermore, we find that an increase in $\mu$ results in a reduced speedup. This occurs because CD computes the combined distribution by subtracting the proposal model’s information from the target model’s. As $\mu$ grows, the gap between these distributions widens, lowering the acceptance rate (Equation 7). Despite this, the speedup ratio remains above 1.00x, confirming that CD-CoS always accelerates, consistent with Corollary 3.7.
@@ 5 Conclusion
This paper introduces Collaborative Decoding via Speculation (CoS), an extension of speculative decoding that accelerates LLM collaborative decoding while maintaining output quality. CoS refines the verification mechanism for direct collaborative sampling and introduces an alternate proposal framework to further boost efficiency. We demonstrate the effectiveness of CoS through both theoretical analysis and empirical validation.
Acknowledgements
This work is supported by the National Science Foundation of China (62206048), the Natural Science Foundation of Jiangsu Province (BK20220819), and the Fundamental Research Funds for the Central Universities (2242024k30035). Additional support was provided by the Big Data Computing Center of Southeast University and the SEU & Ascend Center of Cultivation.
@@ Appendix A Mathematical Proofs
A.1 Correctness of CoS
Assume that in speculative decoding, the proposal distribution is $q(x)$, the target distribution is $p(x)$, and the combined distribution is $r(x)=\mathcal{C}(q(x),p(x))$. Referring to the proof of speculative decoding correctness (Leviathan et al., 2023), we now prove that the sampling method described in  Section 3.2 ensures that the generated tokens align with the combined distribution $r(x)$.
@@ A.1 Correctness of CoS
Assume that the time required to invoke the target model once is $T$, and the time required to invoke a proposal model is $cT$. Therefore, one proposal and one verification together take $(\gamma c+1)T$ time. The time required to generate one token is $\frac{(1-\alpha)(\gamma c+1)}{1-\alpha^{\gamma}}T$. In the vanilla collaborative decoding, the time required to generate one token is $(c+1)T$. Therefore, the improvement factor in total walltime is $\frac{(1-\alpha^{\gamma})(1+c)}{(1-\alpha)(\gamma c+1)}$.
@@ B.1 Alternate Proposal Framework
Algorithm 1 The alternate proposal framework. Propose takes the current token sequence $S$ and a constant $\gamma$ as inputs and generates $\gamma$ tokens $T$. Score feeds a sequence to the current verifier to obtain logits $L$ and a bonus token $t$. Verify examines $T$ to decide whether it should be accepted according to combined logits.
@@
     $L^{\prime}=\mathcal{C}^{\prime}(L)$
@@
B.2 CoS Framework
The CoS framework is a generalization of the alternate proposal framework to scenarios involving more than three models. Figure 7 illustrates our CoS in a three-model scenario, where the models $\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}$ have proposal lengths of $\gamma_{1}=3$, $\gamma_{2}=2$, and $\gamma_{3}=1$, respectively.